OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-2, 3) and is perpendicular to 2x – 9y = -12 please show work!

OpenStudy (anonymous):

y = 3 = -9/2 (x+2)

OpenStudy (anonymous):

-9/2 is the slope y = -9/2x+m 3 = -9/2(-2) +b b = -6 y = -9/2x-6

OpenStudy (anonymous):

y - 3*

OpenStudy (amistre64):

-b/a = slope of perp; 9/2 y = mx -mPx + Py; P(-2, 3) y = 9/2 x -(9(-2)/2)+3 y = 9/2 x +9+3 y = 9/2 x +12 maybe?

OpenStudy (anonymous):

the slope of the given line is -a/b (a = 2, b = -9) the slope of a line perpendicular to a line with slope m is -1/m so the slope of the line perpendicular to the given line is -1/(-a/b) = b/a = -9/2 then substitute this slope in the slope intercept form y = mx+b substitute the point and the slope you found and find b